OpenStudy (anonymous):

limit (1+(1/x^2))x as x approaches infinity

OpenStudy (anonymous):

using L'Hopitals rule

OpenStudy (anonymous):

This?\[ \lim_{x\to\infty}\left(1+\frac{1}{x^2}\right)x \]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I know the answer is infinity but I need to know how to do it with L'Hopital

OpenStudy (anonymous):

So simplify the inner part a bit: \[ \left(1+\frac{1}{x^2}\right)x=\frac{x^2+1}{x} \]

OpenStudy (anonymous):

Divide everything by \(x\): \[ \frac{x+\frac 1x}{1} \]

OpenStudy (anonymous):

so its 2x/1 and then you get infinity

OpenStudy (anonymous):

You don't need to use l'Hospitals rule on this one.

OpenStudy (anonymous):

This goes to infinity and it isn't an indeterminate form.

OpenStudy (anonymous):

yeah I know but I need to do it with L'Hopital

OpenStudy (anonymous):

Okay, well you know how to do derivatives, I don't see the issue.

OpenStudy (anonymous):

okay I just couldnt figure out how to get it in a form where I could do it in the end it was obvious lol. Thanks for the help.